A charged conducting sphere of radius R is embedded in a material that
is infinite in extent and has a permittivity of apsolon E. The external electric field
is uniform and has a magnitude of Eo. The sphere has a net charge of q.
Find the electric field (a) inside and (b) outside the sphere.
1
Expert's answer
2020-08-14T15:57:40-0400

(a) inside the conducting sphere the field is zero ("E=0" )


(b) the field created by the charge on the sphere


"\overrightarrow{E}=k\frac{q}{r^2}\cdot\frac{\overrightarrow{r}}{r}" ("r>R" )


The field outside the sphere


"\overrightarrow{E}_\Sigma=\overrightarrow{E}_0+k\frac{q}{r^2}\cdot\frac{\overrightarrow{r}}{r}"